OpenStudy (anonymous):

Need hep finding the surface area and volume of the sixth stellation of the icosahedron. I am realllly bumped.

OpenStudy (anonymous):

Are you using multi variables and triple integrals to solve this?

OpenStudy (anonymous):

this is what i am dealing with. to tell you the truth, I have no idea what you are talking about :P I was asked for a project to find the surface area and volume. is there a formula for this crazy polyhedron?

OpenStudy (anonymous):

Are you in Calc 3?

OpenStudy (anonymous):

nope, honors geometry. I am a sophmore

OpenStudy (anonymous):

What methods do you use to solve volumes of shapes? What I'm asking is what is the first step?

OpenStudy (anonymous):

like for a cone we use V=1/3 BH , for a hemisphere, we use 4/3 pi r^3. pretty basic pyramid= 1/3 BH

OpenStudy (anonymous):

So you are given the formula of shapes for volume. Does the book give you the formula in this case?

OpenStudy (anonymous):

no it does not. our teacher was suppose to teach us the formula, but he didn't get around to it or I missed it. and i was sure to pay attention. either way, I have no cue where I am suppose to start.

OpenStudy (anonymous):

Sorry I know a different methods but it's too tedious. Wait and ask your teacher tomorrow

OpenStudy (anonymous):

alright, thank you